# Spike Imaging Velocimetry: Dense Motion Estimation of Fluids Using Spike Streams

### v3

@@ Introduction
DIP adopts the same compact and effective convex upsampling strategy as RAFT to recover full-resolution optical flow. However, such upsampling is insufficient for accurately reconstructing the velocity field of fluids. To address this, we propose a Multi-scale Velocity Refinement (MSVR) module to assist in reconstructing the complete velocity field, including fine-scale vortices. As illustrated in Figure 4, we first apply a sequence of 2D convolutions followed by activation functions to process the multi-scale velocity fields generated by MCIO. Information exchange across scales is achieved via cross convolution from coarse to fine levels. The extracted features are then fused and processed by distinct convolutional networks to generate a residual velocity field $\mathbf{Vel}^{\text{res}}$ and a quality map $\mathbf{Q}$. The refined fluid velocity field is given by,
$\mathbf{Vel}^{*}=\mathbf{Vel}^{i=\text{iters}}+\mathbf{Vel}^{\text{res}}\odot\mathbf{Q},$ (10)
$\odot$ denotes element-wise multiplication.
@@
Since SIV outputs a prediction sequence of length $N$ at each iteration, the flow loss is defined as:
$L_{\text{flow}}=\sum_{i=0}^{i=N}w_{i}\cdot\|f_{i}-f_{gt}\|_{1}.$ (11)
We introduce a gradient loss to prevent the generated flow field from becoming overly smooth in regions like shear layers or small vortical structures. It is computed as follows:
$L_{\text{grad}}=\sum_{i=0}^{i=N}w_{i}\cdot(\|\nabla_{\mathbf{x}}f_{i}-\nabla_{\mathbf{x}}f_{gt}\|_{1}+\|\nabla_{\mathbf{y}}f_{i}-\nabla_{\mathbf{y}}f_{gt}\|_{1}).$ (12)
We set $w_{i}=\gamma^{N-i-1}$ and $\gamma=0.8$ for both $L_{\text{flow}}$ and $L_{\text{grad}}$. The total loss can be formulated as follows:
@@
Here, we set $\beta=0.3$.
Experimental Results
In this section, we evaluate our method on the PSSD dataset and compare its performance with other traditional and learning-based methods. Additionally, we conduct a series of ablation experiments to assess the effectiveness of the proposed modules.
@@
SIV is implemented using the PyTorch framework on a single NVIDIA RTX 2080 Ti GPU. We employ the Adam optimizer with a warm-up strategy during training. For each sub-dataset, we train the model for 100 epochs with an initial learning rate of $10^{-4}$, applying an 85% decay every 10 epochs. The number of iterative refinements of GRUs is set to 12. The batch size is set to 4.
Following (Cai et al. 2019b) and (Lagemann et al. 2021), we use endpoint error (EPE) as the primary evaluation metric for comparing different algorithms. Additionally, we incorporate another widely used PIV metric, Root Mean Squared Error (RMSE), in our extensive experiments, as detailed in the supplementary material.
@@
coarse-to-fine HS 6.790 0.509 7.415 1.998 7.683 2.558 9.388 1.954 7.819 1.755
RAFT-PIV-Image 0.377 0.215 2.181 1.252 1.336 0.842 1.875 1.064 1.442 0.843
RAFT-PIV-Spike 0.347 0.208 1.223 0.755 0.857 0.557 0.958 0.581 0.846 0.525
PIV-NetS-Image 0.545 0.301 2.560 1.497 1.574 0.985 2.247 1.320 1.731 1.026
PIV-NetS-Spike 0.553 0.289 1.507 0.874 1.252 0.734 1.382 0.782 1.173 0.670
SCFlow 0.354 0.222 2.232 1.305 1.374 0.901 2.132 1.169 1.522 0.899
HiST-SFlow 0.369 0.224 1.271 0.796 0.964 0.611 1.029 0.637 0.908 0.567
Flowformer-Spike 0.286 0.147 1.069 0.642 0.727 0.437 0.806 0.472 0.722 0.425
SIV (ours) 0.239 0.159 0.896 0.587 0.613 0.422 0.680 0.441 0.607 0.402
Table 1: EPE ($\downarrow$) results on PSSD dataset (Problem 1). The best results are marked in bold.
Comparative Experiments
We compare our method against several open-source baselines, including a classical PIV algorithm (Cai et al. 2018), PIV networks (PIV-NetS (Cai et al. 2019b), RAFT-PIV (Lagemann et al. 2021)), spike-adapted general optical flow networks (e.g., Flowformer with spike input (Huang et al. 2022)), and spike-based optical flow networks (SCFlow (Hu et al. 2022), HiST-SFlow (Zhao et al. 2024)).
@@
coarse-to-fine HS 20.691 8.244 17.164 8.051 17.469 8.170 17.131 7.936 18.114 8.100
RAFT-PIV-Image 0.665 0.368 4.730 2.635 2.320 1.366 4.305 2.368 3.005 1.684
RAFT-PIV-Spike 0.688 0.412 2.410 1.483 1.468 0.959 2.071 1.241 1.659 1.024
PIV-NetS-Image 1.163 0.683 4.866 2.830 2.673 1.643 4.347 2.500 3.262 1.914
PIV-NetS-Spike 0.887 0.540 3.172 1.818 2.392 1.410 3.067 1.744 2.379 1.378
SCFlow 0.694 0.403 4.689 2.613 2.387 1.391 4.326 2.380 3.024 1.696
HiST-SFlow 0.556 0.364 2.212 1.365 1.404 0.998 1.911 1.179 1.521 0.976
Flowformer-Spike 0.533 0.284 1.975 1.179 1.214 0.748 1.668 0.972 1.348 0.796
SIV (ours) 0.370 0.254 1.687 1.080 1.071 0.722 1.449 0.912 1.144 0.742
Table 2: EPE ($\downarrow$) results on PSSD dataset (Problem 2). The best results are marked in bold.
As shown in Table 1, 2, and 3, $\Delta t=21$ and $\Delta t=11$ represent the number of timestamps for the central temporal intervals between two spike streams used for velocity field prediction, respectively. For detailed descriptions, please refer to the supplementary material. Our proposed SIV method achieves the lowest average End-Point Error (EPE) across all evaluated sub-datasets. Compared to baseline approaches, SIV demonstrates superior prediction accuracy in both low-velocity (Problem 1) and high-velocity (Problem 2) optical flow scenarios. Furthermore, in high dynamic range conditions (Problem 3), SIV effectively exploits the inherent advantages of spike-based input modalities to address the fundamental limitations of conventional frame-based imaging systems.
Figure 5 presents the visualized predictions of all methods across three datasets. The top row shows the velocity fields visualized using the standard optical flow color coding (Baker et al. 2011). The bottom row shows the endpoint error (EPE) maps between the predicted and ground-truth flow. Shear layers are often a primary source of turbulence generation and development. SIV achieves the lowest prediction error in both shear layers and small-scale vortices, while avoiding noticeable blurring, distortion, or discontinuities.
@@
coarse-to-fine HS 7.908 1.044 7.924 2.706 8.345 3.130 7.854 2.716 8.008 2.399
RAFT-PIV-Image 0.447 0.244 2.694 1.495 1.447 0.888 2.311 1.271 1.725 0.974
RAFT-PIV-Spike 0.320 0.200 1.203 0.740 0.798 0.530 0.946 0.576 0.817 0.512
PIV-NetS-Image 0.667 0.343 2.786 1.602 1.641 0.956 2.393 1.378 1.872 1.070
PIV-NetS-Spike 0.476 0.247 1.420 0.856 1.157 0.716 1.342 0.774 1.099 0.648
SCFlow 0.462 0.269 1.850 0.837 1.300 0.774 1.972 0.769 1.396 0.662
HiST-SFlow 0.329 0.180 1.179 0.640 0.832 0.452 0.960 0.476 0.825 0.437
Flowformer-Spike 0.269 0.145 1.052 0.620 0.716 0.429 0.813 0.481 0.713 0.419
SIV (ours) 0.234 0.155 0.880 0.573 0.598 0.404 0.677 0.438 0.598 0.392
Table 3: EPE ($\downarrow$) results on PSSD dataset (Problem 3). The best results are marked in bold.
@@
Ablation of modules. We conduct a series of ablation experiments to evaluate the contribution of each component to overall performance, as detailed in Table 4. All the values are the arithmetic mean of the scenes. The best results are marked in bold. These results underscore the effectiveness of the DPHT, GE, MSVR and $L_{\text{grad}}$. In Table 4, the comparison between experiments (A) and (B) demonstrates the effectiveness of DPHT. Both the comparison between (B) and (C) and that between (D) and (E) demonstrate the effectiveness of the Graph Encoder (GE). Both the comparison between (B) and (D) and that between (C) and (E) highlight the contribution of MSVR.
Ablation of $\beta$. Table 5 summarizes the ablation study on the weighting coefficient $\beta$ of the proposed loss function. The results indicate that setting $\beta=0.3$ yields the best performance for the current model and dataset. Additional ablation studies are provided in the supplementary material.
@@
Table 4: Ablation study of proposed modules, showing results for EPE ($\downarrow$). $\checkmark$ indicates module presence, $\times$ indicates absence. Setting (G) is our final model.
@@ Conclusion
We explore a novel approach to fluid motion estimation using spike cameras and propose a specialized network architecture, termed Spike Imaging Velocimetry (SIV), designed to solve PIV tasks while adapting to the scenario-specific characteristics of fluid scenes. We propose a Detail-Preserving Hierarchical Transform module to preserve particle signal information during the pre-processing of spike streams. We design a Graph Encoder to extract contextual features from highly unstructured and turbulent flow fields. We develop a Multi-scale Velocity Refinement module to refine the velocity field. We also introduce a challenging and realistic PIV dataset, named Particle Scenes with Spike and Displacement (PSSD). Experimental results demonstrate that our method achieves state-of-the-art performance on the PSSD dataset and accurately estimates velocities in small-scale yet critical flow structures such as shear layers and vortices. In summary, this study demonstrates the potential of spike cameras in fluid velocity measurement and offers new perspectives and directions for PIV network design.

### v4

@@ Introduction
DIP adopts the same compact and effective convex upsampling strategy as RAFT to recover full-resolution optical flow. However, such upsampling is insufficient for accurately reconstructing the velocity field of fluids. To address this, we propose a Multi-scale Velocity Refinement (MSVR) module to assist in reconstructing the complete velocity field, including fine-scale vortices. As illustrated in Figure 5, we first apply a sequence of 2D convolutions followed by activation functions to process the multi-scale velocity fields generated by MSIO. Information exchange across scales is achieved via cross convolution from coarse to fine levels. The extracted features are then fused and processed by distinct convolutional networks to generate a residual velocity field $\mathbf{u}_{\text{res}}$ and a quality map $\mathbf{Q}$. The refined fluid velocity field is given by,
$\mathbf{u}_{\text{ref}}=\mathbf{u}_{N}+\mathbf{u}_{\text{res}}\odot\mathbf{Q},$ (10)
where $\odot$ denotes element-wise multiplication.
@@
The loss function for the proposed network is composed of two parts: the flow loss and the gradient loss. Suppose the recurrent optimizer of the network has $N$ iterations, and the estimated flow fields of each iteration are $\{\mathbf{u}_{1},\dots,\mathbf{u}_{N},\mathbf{u}_{N+1}\}$, $\mathbf{u}_{N+1}=\mathbf{u}_{\text{ref}}$. The flow loss is defined as:
$L_{\text{flow}}=\sum_{i=1}^{N+1}\gamma^{N+1-i}\cdot\|\mathbf{u}_{i}-\mathbf{u}_{\text{gt}}\|_{1},$ (11)
$\mathbf{u}_{\text{gt}}$ is the ground truth of velocity field. To preserve small vortical structures, we introduce a gradient loss:
$L_{\text{grad}}$ $=\sum_{i=1}^{N+1}\gamma^{N+1-i}\cdot(\|\nabla_{\mathbf{x}}\mathbf{u}_{i}-\nabla_{\mathbf{x}}\mathbf{u}_{\text{gt}}\|_{1}$ (12)
$+\|\nabla_{\mathbf{y}}\mathbf{u}_{i}-\nabla_{\mathbf{y}}\mathbf{u}_{\text{gt}}\|_{1}).$
The total loss is:
@@
$\gamma$ is set to 0.8, and $\beta$ is set to 0.3.
Experiments
We evaluate our method on the PSSD dataset and compare its performance with both traditional and learning-based approaches. Additionally, we conduct a series of ablation studies to assess the effectiveness of the proposed modules.
@@
SIV is implemented using the PyTorch framework on a single NVIDIA RTX 2080 Ti GPU. We use the Adam optimizer with a warm-up strategy during training. For each sub-dataset, the model is trained for 100 epochs with a batch size of 4 and an initial learning rate of $10^{-4}$, applying an 85% decay every 10 epochs. The number of iterative refinements for the GRUs is set to 12. Following (pivnets) and (raftpiv), we use endpoint error (EPE) as the primary evaluation metric for comparing different algorithms. The EPE is the spatially mean value of Euclidean distance between the predicted flow $\mathbf{u}_{\text{pred}}$ and its ground truth $\mathbf{u}_{\text{gt}}$.
@@
(A) coarse-to-fine HS 6.790 0.509 7.415 1.998 7.683 2.558 9.388 1.954 7.819 1.755
(B) RAFT-PIV-Image 0.377 0.215 2.181 1.252 1.336 0.842 1.875 1.064 1.442 0.843
(C) RAFT-PIV-Spike 0.347 0.208 1.223 0.755 0.857 0.557 0.958 0.581 0.846 0.525
(D) PIV-NetS-Image 0.545 0.301 2.560 1.497 1.574 0.985 2.247 1.320 1.731 1.026
(E) PIV-NetS-Spike 0.553 0.289 1.507 0.874 1.252 0.734 1.382 0.782 1.173 0.670
(F) SCFlow 0.354 0.222 2.232 1.305 1.374 0.901 2.132 1.169 1.522 0.899
(G) HiST-SFlow 0.369 0.224 1.271 0.796 0.964 0.611 1.029 0.637 0.908 0.567
(H) Flowformer-Spike 0.286 0.147 1.069 0.642 0.727 0.437 0.806 0.472 0.722 0.425
(I) SIV (ours) 0.239 0.159 0.896 0.587 0.613 0.422 0.680 0.441 0.607 0.402
Table 1: EPE ($\downarrow$) results on PSSD dataset - Problem 1. The best results are marked in bold.
@@
(A) coarse-to-fine HS 20.691 8.244 17.164 8.051 17.469 8.170 17.131 7.936 18.114 8.100
(B) RAFT-PIV-Image 0.665 0.368 4.730 2.635 2.320 1.366 4.305 2.368 3.005 1.684
(C) RAFT-PIV-Spike 0.688 0.412 2.410 1.483 1.468 0.959 2.071 1.241 1.659 1.024
(D) PIV-NetS-Image 1.163 0.683 4.866 2.830 2.673 1.643 4.347 2.500 3.262 1.914
(E) PIV-NetS-Spike 0.887 0.540 3.172 1.818 2.392 1.410 3.067 1.744 2.379 1.378
(F) SCFlow 0.694 0.403 4.689 2.613 2.387 1.391 4.326 2.380 3.024 1.696
(G) HiST-SFlow 0.556 0.364 2.212 1.365 1.404 0.998 1.911 1.179 1.521 0.976
(H) Flowformer-Spike 0.533 0.284 1.975 1.179 1.214 0.748 1.668 0.972 1.348 0.796
(I) SIV (ours) 0.370 0.254 1.687 1.080 1.071 0.722 1.449 0.912 1.144 0.742
Table 2: EPE ($\downarrow$) results on PSSD dataset - Problem 2. The best results are marked in bold.
@@
(A) coarse-to-fine HS 7.908 1.044 7.924 2.706 8.345 3.130 7.854 2.716 8.008 2.399
(B) RAFT-PIV-Image 0.447 0.244 2.694 1.495 1.447 0.888 2.311 1.271 1.725 0.974
(C) RAFT-PIV-Spike 0.320 0.200 1.203 0.740 0.798 0.530 0.946 0.576 0.817 0.512
(D) PIV-NetS-Image 0.667 0.343 2.786 1.602 1.641 0.956 2.393 1.378 1.872 1.070
(E) PIV-NetS-Spike 0.476 0.247 1.420 0.856 1.157 0.716 1.342 0.774 1.099 0.648
(F) SCFlow 0.462 0.269 1.850 0.837 1.300 0.774 1.972 0.769 1.396 0.662
(G) HiST-SFlow 0.329 0.180 1.179 0.640 0.832 0.452 0.960 0.476 0.825 0.437
(H) Flowformer-Spike 0.269 0.145 1.052 0.620 0.716 0.429 0.813 0.481 0.713 0.419
(I) SIV (ours) 0.234 0.155 0.880 0.573 0.598 0.404 0.677 0.438 0.598 0.392
Table 3: EPE ($\downarrow$) results on PSSD dataset - Problem 3. The best results are marked in bold.
Comparative Experiments
We compare our method with several open-source baselines, including classical PIV algorithms (cai2018motion), image-based PIV networks (PIV-NetS (pivnets), RAFT-PIV (raftpiv)), spike-adapted optical flow networks (e.g., Flowformer with spike input (flowformer)), and spike-based optical flow networks (SCFlow (hu2022optical), HiST-SFlow (hist)). All networks trained on spikes (incl. (C)(E)(F)(G)(H)(I) in Table 1, 2, and 3) use voxel-grid inputs to ensure a fair evaluation. All image baselines (incl. (A)(B)(D) in Table 1, 2, and 3) use sharp frames containing only artificially added noise (as shown in Figure 2, not spike reconstructions) simulating high-speed cameras to highlight the differences between image and spike inputs. The image is temporally aligned with the intermediate moments of the spike stream.
As shown in Tables 1, 2, and 3, $\Delta t=21$ and $\Delta t=11$ represent the time intervals between two spike streams used for velocity field prediction. Our SIV method achieves the lowest average EPE across all sub-datasets. Compared to baseline approaches, SIV demonstrates superior prediction accuracy in both low-velocity (Problem 1) and high-velocity (Problem 2) scenarios. Additionally, in high dynamic range conditions (Problem 3), SIV effectively leverages the advantages of spike-based input modalities, addressing the fundamental limitations of conventional frame-based systems.
Figure 6 visualizes the predictions on three problems. The top row visualizes the velocity fields using the standard optical flow color coding (colorcoding), and the bottom row shows the EPE. Since shear layers are crucial for turbulence generation and development, it is noteworthy that SIV attains the lowest prediction error in both shear layers and small-scale vortices, while avoiding noticeable blurring, distortion, or discontinuities.
@@
Ablation of modules. All values in Table 4 and 5 are the arithmetic mean across scenes ($\Delta t=21$), with the best results highlighted in bold. The results emphasize the effectiveness of DPHT, GE, MSVR, and $L_{\text{grad}}$. In Table 4, comparisons between experiments (A) and (B) show the effectiveness of DPHT, while comparisons between (B) and (C), and (D) and (E), validate the contribution of the Graph Encoder (GE). Additionally, comparisons between (B) and (D), and (C) and (E), highlight the impact of MSVR.
Ablation of $\beta$. Table 5 summarizes the ablation study on the weighting coefficient $\beta$ in the proposed loss function. The results show that $\beta=0.3$ yields the best performance for the current model and dataset.
@@
Table 4: Ablation study of proposed modules, showing results for EPE ($\downarrow$). $\checkmark$ indicates module presence, $\times$ indicates absence. Setting (E) is our final model.
@@ Conclusion
We propose a novel approach for fluid motion estimation using spike cameras and introduce a specialized network architecture, Spike Imaging Velocimetry (SIV), designed to address PIV tasks while adapting to the specific characteristics of fluid scenes. We develop the Detail-Preserving Hierarchical Transform (DPHT) module to preserve particle signal information during spike stream preprocessing, a Graph Encoder (GE) to extract contextual features from highly unstructured and turbulent flows, and a Multi-scale Velocity Refinement (MSVR) module to refine the velocity field. We also introduce the challenging and realistic Particle Scenes with Spike and Displacement (PSSD) dataset. Experimental results demonstrate that SIV achieves state-of-the-art performance on the PSSD dataset, accurately estimating velocities in small-scale yet critical flow structures such as shear layers and vortices. In summary, this study highlights the potential of spike cameras for fluid velocity measurement and offers new perspectives and directions for PIV network design.
This paper aims to demonstrate that spike cameras are a promising tool for PIV estimation. However, given the complexity of fluid flows, there remains substantial room for improvement: (i) although we have generated a large dataset, its scale and diversity should be further increased; (ii) noise should be modeled more comprehensively to achieve stronger performance on experimental data; and (iii) improving inference efficiency to enable real-time prediction would benefit both laboratory and industrial applications.
Acknowledgements
This work is supported by the Postgraduate Research & Practice Innovation Program of Jiangsu Province (No. KYCX25_0173), the National Natural Science Foundation of China under Grants 62371006, 62302016, U24B20140, the Jiangsu Association for Science and Technology (JAST) Young Elite Scientists Sponsorship Program (No. JSTJ-2025-730), and the Beijing Natural Science Foundation (No. 3242008).